Consider a 15.60g sample of 131I, which has a half-life of 8.02 days.

a. How much of the sample (in grams) will remain after 16.04 days?

b. How much of the sample will remain after 16.04 hours?
Answer #1

a)

half - life = 8.02 days

rate constant k = 0.693 / t1/2

                        = 0.693 / 8.02

                        = 0.0864 day-1

rate constant k = 1/t ln (Ao / At)

0.0864 = 1/ 16.04 ln (15.60 / At)

At = 3.90 g

mass of sample remain = 3.901 g

b)

time = 16.04 hours = 0.6683 days

rate constant k = 1/t ln (Ao / At)

0.0864 = 1/ 0.668 ln (15.60 / At)

At = 14.72 g

mass of sample remain = 14.72 g